Solve the following equations by Elimination method, Substitution method and Graphing method.


1. 2π‘₯ + 3𝑦 = 8


5π‘₯ βˆ’ 3𝑦 = βˆ’1


2. 2π‘₯ +5y = 19


x βˆ’ 2y = βˆ’4


3. 𝑦 = 5 βˆ’ 2π‘₯


4π‘₯ + 3𝑦 = 13


4. 𝑦 = 2π‘₯ βˆ’ 3


𝑦 = βˆ’ 2/3π‘₯ + 5


1
Expert's answer
2022-03-15T09:33:14-0400

Substitution method:

1.x=(8-3y)/2

(40-15y)/2-3y=-1

-10.5y=-21

y=2

x=1

2.x=-4+2y

-8+4y+5y=19

9y=27

y=3

x=2

3.4x+15-6x=13

2x=2

x=1

y=3

4.2x-3=-2/3x+5

8/3x=8

x=3

y=3

Elemination method

1. Summarize equalations:

7x=7

x=1

3y=8-2=6

y=2

2. Multiply second equalation on -2 and summarize: 9y=27

y=3

x=-4+6=2

3.Multiply first equalation on -3 and summarize: 4x=-15+6x+13

2x=2

x=1

y=3

4.Multiply second equalation on -1 and summarize:0=8/3x-8

8/3x=8

x=3

y=6-3=3

Graphic method